how do you calculate x when -1/x = -0.125
Found 2 solutions by macston, Alan3354:
Answer by macston(5194)

-1/x=-0.125 Multiply each side by x
-1=-0.125x Divide each side by -0.125
-1/-0.125=x
8=x ANSWER: x equals 8

Answer by Alan3354(69443)

how do you calculate x when -1/x = -0.125
------
Multiply by -1
Invert.